# Retrieval-Augmented Generation for AI-Generated Content: A Survey

### v3

@@ II-A Overview
II-B Generator
(a) General architectures of several generators.
@@ II-B Generator
II-B1 Transformer Model
Transformer models are one of the best performing models in the field of Natural Language Processing (NLP), consisting of self-attention mechanisms, feed-forward networks, layer normalization modules, and residual networks [62]. As shown in Fig. 5(a), the input of the transformer is mapped to a tensor $x_{in}$ with a shape of (b, s, h) after the tokenization process and embedding model, where b represents batch size, s represents sequence length and h represents hidden dimension. Next, the position encoding will be sent to the self attention layer along with this tensor. The input $x_{in}$ and the output $x_{out}$ of the self-attention module will be connected by the residual network and the layer normalization module. Finally, the output of the ”Add & Norm” module $x_{out}$ will be sent to the feed forward network.
The entire process can be defined as follows:
$Q=x_{in}*W_{q}+b_{q}$
$K=x_{in}*W_{k}+b_{k}$
$V=x_{in}*W_{v}+b_{v}$
$x_{out}=LayerNorm1(Softmax(\frac{Q*K^{T}}{\sqrt{h}})*V*W_{o}+b_{o})+x_{in}$
$y=LayerNorm2((x_{out}*W_{1}+b_{1})*W_{2}+b_{2})+x_{out}$
It should be noted that $w_{q},w_{k},w_{v},w_{o}$ are learnable tensors with shape (h, h); $b_{q},b_{k},b_{v},b_{o}$ are a learnable tensors with shape (h,).
II-B2 LSTM
@@
Long Short-Term Memory (LSTM) [63] is a special Recurrent Neural Network (RNN) model that overcomes the exploding/vanishing gradient problems of RNN in processing long-term dependency information by introducing cell state and gate mechanisms. The LSTM model consists of three gates: Input Gate, Forget Gate, and Output Gate. These gates update the cell state by controlling the information flow, enabling the model to remember long-term dependent information. Cell State is the core module of the LSTM model which can memorize and maintain information. The Input Gate decides which input data should be retained in the cell state. Forget Gate determines which cell state information should be discarded to avoid excessive memory. Output Gate determines how the information in the cell state affects the current output. The flow of data and the collaborative work process between components are shown in the Fig. 5(a).
The entire process can be defined as follows:
$f=sigmoid(W^{f}*x^{t}+U^{f}*y^{t-1}+b^{f})$
$z=tanh(W^{z}*x^{t}+U^{z}*y^{t-1}+b^{z})$
$i=sigmoid(W^{i}*x^{t}+U^{i}*y^{t-1}+b^{i})$
$o=sigmoid(W^{o}*x^{t}+U^{o}*y^{t-1}+b^{o})$
$c^{t}=z\odot i+f\odot c^{t-1}$
$y^{t}=o\odot tanh(c^{t})$
II-B3 Diffusion Model
@@
Diffusion models [64] are a family of deep generative models that can create realistic and diverse samples of data [65, 66, 67, 68, 69, 70, 71, 72], such as images [73, 74, 75, 76, 77, 78, 79], texts [80, 81, 82, 83], videos [84, 85, 86, 87, 88], and molecules [89, 90, 91, 92, 93]. As shown in Fig. 5(a), diffusion models work by gradually adding noise to data until it becomes random, then reversing the process to generate new data from noise. This process is based on probabilistic modeling and neural networks. Diffusion models mainly have three equivalent formulations: denoising diffusion probabilistic models [65, 66, 67], score-based generative models [68, 69], and stochastic differential equations [70, 71], with following improvements like DDIM [94], Rectified Flow [95], Consistency Model [96] and RPG-DiffusionMaster [79].
Especially, let $x_{0}$ be a random variable that follows the data distribution $q(x_{0})$, and let $x_{t}$ be a random variable that follows the distribution $q(x_{t}|x_{0})$ after adding noise at time step $t$. Then, DDPM can be formulated as follows:
  • Forward Process The forward process perturbs data with a sequence of Gaussian noise injections, transforming the data distribution $q(x_{0})$ into a simple prior distribution $q(x_{T})\approx N(0,I)$. The transition kernel at each time step is given by
    $q(x_{t}|x_{t-1})=N(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I),$
    where $\beta_{t}\in(0,1)$ is a hyperparameter. The marginal distribution of $x_{t}$ conditioned on $x_{0}$ is
    $q(x_{t}|x_{0})=N(x_{t};\sqrt{\bar{\alpha}_{t}}x_{0},(1-\bar{\alpha}_{t})I),$
    where $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}=\prod_{s=0}^{t}\alpha_{s}$.
  • Reverse Process The reverse process generates new data samples by reversing the forward process with a learnable Markov chain. The prior distribution is $p(x_{T})=N(x_{T};0,I)$ and the transition kernel is
    $p_{\theta}(x_{t-1}|x_{t})=N(x_{t-1};\mu_{\theta}(x_{t},t),\Sigma_{\theta}(x_{t},t)),$
    where $\theta$ denotes model parameters, and $\mu_{\theta}(x_{t},t)$ and $\Sigma_{\theta}(x_{t},t)$ are parameterized by deep neural networks. The reverse process starts from sampling $x_{T}\sim p(x_{T})$ and iteratively samples $x_{t-1}\sim p_{\theta}(x_{t-1}|x_{t})$ until $t=0$.
  • Model Training For each sample $x_{0}\sim q(x_{0})$, the model training objective is to maximizing the variational lower bound (VLB) of the log-likelihood of the data $x_{0}$. The simplified form $\mathcal{L}_{\text{VLB}}(x_{0})$ is given by
    $\mathbb{E}_{q(x_{1:T}|x_{0})}\left[-\log p(x_{T})-\sum_{t=1}^{T}\log\frac{p_{\theta}(x_{t-1}|x_{t})}{q(x_{t}|x_{t-1})}\right]$
    With simplication and reparameterization trick, the overall objective $\mathbb{E}_{q(x_{0})}\left[\mathcal{L}_{\text{VLB}(x_{0})}\right]$ can be simplified into the final form
    $\mathbb{E}_{t\sim U[1,T],x_{0}\sim q(x_{0}),\epsilon\sim\mathcal{N}(0,I)}\left[\lambda(t)\|\epsilon-\epsilon_{\theta}(x_{t},t)\|\right]$
    where $\lambda(t)$ is a positive weighting function, $U[1,T]$ is a uniform distribution over the set $\{1,2,\dots,T\}$, and $\epsilon_{\theta}$ is a deep neural network with parameter $\theta$ that predicts the noise vector $\epsilon$ given $x_{t}$ and $t$. Note that, the overall objective is also equivalent to matching the joint distribution of the reverse process $p_{\theta}(x_{0},x_{1},\dots,x_{T})$ to that of the forward process $q(x_{0},x_{1},\dots,x_{T})$ by minimizing the KL divergence between them.
II-B4 GAN
Generative Adversarial Networks (GANs) [14] are highly anticipated deep learning models with amazing capabilities which can simulate and generate realistic images, audio, and other data. Due to its outstanding performance, GANs have achieved significant achievements in various fields [97]. The design inspiration of GANs comes from the zero-sum game in game theory.
@@ II-B4 GAN
As shown in Fig. 5(a), a typical GAN consists of two main components: a generator and a discriminator. These two parts compete with each other through adversarial learning, allowing the generator to continuously improve its ability to generate realistic samples, while the discriminator continuously improves its ability to distinguish between true and false samples.
II-C Retriever
@@
Retrieval is to identify and obtain relevant information given an information need. Specifically, let’s consider information resources that can be conceptualized as a key-value store $\{(k_{i},v_{i})\}_{i=1}^{N}$, where each key $k_{i}$ corresponds to a value $v_{i}$ ($k_{i}$ and $v_{i}$ can be identical). Given a query $q$, the objective is to search for the top-$k$ most similar keys using a similarity function $s$, and obtain the paired values. Based on different similarity functions, existing retrieval methods can be categorized into sparse retrieval, dense retrieval, and others. For widely used sparse and dense retrieval, the whole process can be divided into two distinct phases: in the first phase, each object is encoded into a specific representation; and in the second phase, an index is constructed to organize the data source for efficient search.
II-C1 Sparse Retriever
@@
Sparse retrieval methods are widely used in document retrieval, where the keys are actually documents to be searched (values are the same documents in this scenario). These methods leverage term matching metrics such as TF-IDF [98], query likelihood [99], and BM25 [19], which analyze word statistics from texts and construct inverted indices for efficient searching. Among them, BM25 is a hard-to-beat baseline in industrial-scale web search. For a query $q$ containing keywords $\{q_{i}\}_{i=1}^{n}$, the BM25 score of a document $D$ is:
$s(q,D)=\sum_{i=1}^{n}I\!D\!F(q_{i})\cdot\frac{f(q_{i},D)\cdot(a+1)}{f(q_{i},D)+a\cdot(1-b+b\cdot\frac{|D|}{avgdl})}$
where IDFI\!D\!F is the inverse document frequency weight, $f(q_{i},D)$ is the number of times that $q_{i}$ occurs in the document $D$, $|D|$ is the length of $D$, avgdlavgdl is the average document length in the corpus collection, $a$ and $b$ are tunable parameters.
IDFI\!D\!F is computed as:
$I\!D\!F(q_{i})=\ln(\frac{N-n(q_{i})+0.5}{n(q_{i})+0.5}+1)$
where $N$ is the number of documents, and $n(q_{i})$ is the number of documents containing $q_{i}$. IDFI\!D\!F score is also used in TF-IDF.
@@
II-C2 Dense Retriever
Unlike sparse retrieval, dense retrieval methods represent queries and keys using dense embedding vectors, and build approximate nearest neighbor (ANN) index to speed up the search. This can be applied to all modalities. For text data, recent advancements in pre-trained models, including BERT [15] and RoBERTa [100], have been employed to encode queries and keys individually [20, 101, 102, 103, 104]. Similar to text, models have been proposed to encode code data [25, 105, 106], audio data [107, 26], image data [24, 108], video data [109, 110]. etc. The similarity score between dense representations are usually computed with metrics such as cosine, inner product, L2-distance.
During training, dense retrieval usually follows a contrastive learning paradigm, making positive samples more similar and negative samples less similar. Several hard negative techniques [101, 111] have been proposed to further improve the model quality. During inference, approximate nearest neighbor (ANN) methods are applied for efficient searching. Various indices are developed to serve ANN search, such as tree [112, 113], locality sensitive hashing [114], neighbor graph index (e.g., HNSW [115], DiskANN [116], HMANN [117]), and the combination of graph index and inverted index (e.g., SPANN [22]).
II-C3 Others
@@ III-A1 Query-based RAG
Query-based RAG, originated from the idea of prompt augmentation, integrates the user’s query with insights from information fetched during the retrieval process, directly into the initial stage of the language model’s input. This paradigm stands as a widely adopted approach within the applications of RAG. After retrieval, the retrieved content will be merged with the original user query to create a composite input sequence which is subsequently fed into the generator for response generation. Query-based RAG has been widely applied across various modalities.
For Text Generation, REALM [33] employs a dual-BERT framework to streamline knowledge retrieval and integration, marrying pre-trained models with knowledge extractors. The initial BERT module processes the input question alongside documents to facilitate retrieval, utilizing MIPS for selecting the top-k documents with the highest probability and periodically updating the index. The document snippets obtained are then integrated with the query, feeding into the second BERT module to produce multiple outputs that are aggregated into a singular, comprehensive response. Lewis et al. [34] synergized pre-trained language models with knowledge retrieval mechanisms, leveraging DPR and BART structures to accomplish retrieval-augmented generation tasks. DPR serves as the retrieval component, sourcing pertinent information from vast document databases, while BART uses this information for text generation. RAG-Token and RAG-Sequence differ in their retrieval timings, with the former retrieving information at each token generation and the latter conducting a single retrieval for the entire sequence. SELF-RAG [126] enhances the accuracy and relevance of responses by integrating a retrieval and critique strategy. Initially, the model employs a retriever to search for information paragraphs closely related to the input question. Subsequently, the critique model evaluates these paragraphs to determine their relevance and level of support of the retrieved text, assessing their impact on the generation of responses. Finally, the generator model constructs responses based on this information and evaluates the quality of these responses through critique marks. In addition to being compatible with local generators, Query-based RAG is also applicable to scenarios that use LLM through API calls. REPLUG [127] illustrates this methodology by treating the language model as a “black box”, utilizing Contriever to seamlessly incorporate relevant external documents into the query. REPLUG LSR, a variant with LM-Supervised Retrieval, further refines this process by optimizing retrieval through language model-supervised insights, aiming to reduce perplexity scores and improve model performance by enriching its contextual understanding. In-Context RALM [128] uses BM25 for document retrieval and trains a predictive reranker to reorder and integrate the top-ranked documents.
In contemporary research on other modalities, augmenting inputs with retrieved contents (which are not limited to texts) has proven highly effective in enhancing the performance of various tasks. This strategy is applicable across several critical domains, including code generation, audio generation, and Knowledge Base Question Answering (KBQA).
For Text-to-Code task, APICoder [129] and DocPrompting [42] demonstrate how effectively integrating retrieved information into language models can improve the accuracy and relevance of generated code. In Automatic Program Repair task, CEDAR [130] and InferFix [131] utilize retrieved code snippets to aid the repair process, enhancing the model’s understanding and application of repair strategies by combining them with the original input. For Code Completion task, ReACC [132] employs a prompting mechanism, leveraging retrieved code snippets as part of the new input to increase the accuracy and efficiency of code completion.
@@
Recent researches in Knowledge Base Question Answering (KBQA) has also shown significant effects of combining retrieval and language models. For instance, Uni-Parser [133], RNG-KBQA [123], and ECBRF [134] effectively improve the performance and accuracy of QA systems by merging queries and retrieved information into prompts. BLLM augmentation [135] represents an innovative attempt at zero-shot KBQA using black-box large language models. This method, by directly integrating retrieved information into the model input without the need for additional sample training, demonstrates the great potential of combining retrieval and language models to enhance the model’s generalization ability in understanding and answering unseen questions.
@@
In the AI-for-Science domain, Chat-Orthopedist [136] provides support for shared decision-making among adolescents with idiopathic scoliosis. It enhances the application effectiveness and information accuracy of LLMs by integrating retrieved information into the prompts of the model.
@@ III-A2 Latent Representation-based RAG
The Fusion-in-Decoder (FiD) [35] technique leverages both BM25 and DPR for sourcing supportive paragraphs. It concatenates each retrieved paragraph and its title with the query, processing them individually through the encoder. FiD reduces computational complexity and efficiently utilizes relevant information to generate answers by fusing information from multiple retrieved paragraphs in the decoder, rather than processing each paragraph in the encoder. The application of Fusion-in-Decoder methodologies transcends the realm of textual content processing, demonstrating substantial potential and adaptability in processing code, structured knowledge, and diverse multimodal datasets. Specifically within the code-related domain, technologies such as EDITSUM [138], BASHEXPLAINER [139], and RetrieveNEdit [140] adopt the FiD approach, facilitating integration through encoder-processed fusion. Re2Com [141], and RACE [142] , among other methods, also feature the design of multiple encoders for different types of inputs.
In the field of Science, RetMol [55] also employs the Fusion-in-Decoder strategy, integrating information at the decoder stage to enhance the relevance and quality of the generated molecular structures.
In the field of Knowledge Base Question Answering (KBQA), the FiD method has been widely adopted, demonstrating significant effectiveness. UniK-QA [143], DECAF [144], SKP [145], KD-CoT [146], and ReSKGC [147] have effectively enhanced the performance of QA systems through the application of Fusion-in-Decoder technology. This illustrates that by integrating RAG for KBQA, the
Retro [36] pioneers the integration of retrieved text via “Chunked Cross-Attention” a novel mechanism that segments the input sequence into discrete chunks. Each chunk independently executes cross-attention operations, thereby mitigating computational burdens. This technique enables the model to selectively retrieve and assimilate distinct documents for varied sequence segments, fostering dynamic retrieval throughout the generation process. This enhances the model’s adaptability and enriches the contextual backdrop of generated content. In the domain of image generation, cross-attention mechanisms have been widely adopted within RAG frameworks. Methods such as Re-imagen [148], KNN-Diffusion [149], RDM [150] and LAION-RDM & ImageNet-RDM [151] utilize cross-attention to integrate multiple retrieval results, effectively enhancing the overall performance of the models.
In addition, there are also some other novel structures worth our attention, Li [152] introduced the ACM, a text-image affine combination module, which notably does not employ any form of attention mechanism. Memorizing Transformers [31] revolutionize long document processing through the integration of a kNN-augmented attention mechanism within a Transformer layer. This innovation triggers a kNN search amidst input sequence processing, fetching data based on similarities between the sequence and stored key-value pairs, thereby elevating performance without necessitating complete retraining. This approach not only bolsters processing efficiency but also broadens the model’s memory span, enabling self-retrieval from its generated outputs and fine-tuning for extensive knowledge bases or code repositories. Unlimiformer [153], by embedding a k-nearest neighbors (kNN) index within a pre-trained encoder-decoder transformer framework, pioneers handling inputs of indefinite length. Storing hidden states of input tokens in the kNN index allows for the efficient retrieval of highly relevant tokens during decoding. This innovation extends the model’s capacity to manage prolonged sequences. Kuratov et al. [154] integrated Transformer with RNN, utilizing the model’s intermediate output as the content for retrieval. This process was executed at each layer of the Transformer, thereby significantly extending the text window’s length and effectively mitigating the issue of “Lost in the Middle”[155]. Diverging from prior methods for knowledge, EaE [156] empowers language models to internalize explicit entity knowledge. EaE introduces an entity-specific parameterization, optimizing inference efficacy through an entity memory layer embedded within the transformer architecture. This layer directly acquires entity representations from textual data, utilizing a sparse retrieval strategy to fetch the nearest entities based on their embeddings, thus refining the model’s comprehension through a calculated aggregation of entity-specific information. on this basis , TOME [157] shifts the focus towards comprehensive mention encodings, prioritizing the granularity of mention over mere entity representations. It meticulously creates a database that stores key and value encodings along with entity IDs, enabling the retrieval of much more fine-grained information. TOME integrates an initial transformer block to process input texts, followed by TOME blocks with memory attention layers, facilitating the synthesis of multifaceted information sources and enhancing inferential reasoning capabilities, even for unencountered entities.
@@
For video captioning, R-ConvED [48] uses a convolutional encoder-decoder network to process retrieved video-sentence pairs with an attention mechanism, generating hidden states to produce captions. CARE [160] introduces a concept detector to produce concept probabilities, and incorporates concept representations into a hybrid attention mechanism. EgoInstructor [49] employs gated-cross attention to integrate textual inputs with encoded video features, enhancing the relevance and coherence of the generated captions for egocentric video content.
Finally, Latent Representation-based RAG is adaptable to various modalities and tasks. It obtains the hidden states of retrieved data, enabling seamless integration between retrievers and generators. However, additional training is required to align the latent space. Within this paradigm, we have the flexibility to design intricate and novel algorithms that effectively combine the information from retrieved contents.

### v4

@@ II-A Overview
Figure 2: General architectures of several generators.
@@
II-A1 Transformer Model
Transformer models are one of the best performing models in the field of Natural Language Processing (NLP), consisting of self-attention mechanisms, feed-forward networks, layer normalization modules, and residual networks [62]. As illustrated in Figure 2, the final output sequence is produced by executing vocabulary classification at each generative step on a sequence of latent representations derived from tokenization and embedding.
II-A2 LSTM
@@
As shown in Fig. 2, Long Short-Term Memory (LSTM) [63] is a special form of Recurrent Neural Network (RNN) model. It addresses the exploding/vanishing gradient problems when processing long-term dependency information by introducing cell state and gate mechanisms. The model consists of three gates (Input Gate, Forget Gate, and Output Gate), which serve as filters of information, and a core module Cell State, which can memorize and maintain information. It uses the same vocabulary classification method as transformer models to autoregressively generate outputs.
II-A3 Diffusion Model
@@
Diffusion models are a family of deep generative models that can create realistic and diverse samples of data (including images, texts, videos, molecules, etc.) [64]. As shown in Fig. 2, diffusion models work by gradually adding noise to data until it becomes random, then reversing the process to generate new data from noise. This process is based on probabilistic modeling and neural networks.
II-A4 GAN
@@
Generative Adversarial Networks (GANs) [14] are highly anticipated deep learning models which can simulate and generate realistic images, audio, and other data [65]. As shown in Fig. 2, a typical GAN consists of two main components: a generator and a discriminator. These two parts compete with each other through adversarial learning, allowing the generator to continuously improve its ability to generate realistic samples, while the discriminator continuously improves its ability to distinguish between true and false samples.
II-B Retriever
@@
Retrieval is to identify and obtain relevant information given an information need. Specifically, let’s consider information resources that can be conceptualized as a key-value store, where each key corresponds to a value (keys and values can be identical). Given a query, the objective is to search for the top-$k$ most similar keys using a similarity function, and obtain the paired values. Based on different similarity functions, existing retrieval methods can be categorized into sparse retrieval, dense retrieval, and others. In widely used sparse and dense retrieval, the entire process can be divided into two distinct phases: (i) each object is first encoded into a specific representation; and then (ii) an index is constructed to organize the data source for efficient search.
II-B1 Sparse Retriever
@@
Sparse retrieval methods are commonly used in document retrieval, where the keys/values represent the documents to be searched. These methods leverage term matching metrics such as TF-IDF [66], query likelihood [67], and BM25 [19], which analyze word statistics from texts and construct inverted indices for efficient searching. Essentially, BM25 is a strong baseline in large-scale web search, integrating inverse document frequency weights, query token occurrences, and other pertinent metrics.
@@
II-B2 Dense Retriever
Unlike sparse retrieval, dense retrieval methods represent queries and keys using dense embedding vectors, and build approximate nearest neighbor (ANN) index to speed up the search. This can be applied to all modalities. For text data, recent advancements in pre-trained models (such as BERT [15]) have been employed encode queries and keys individually [20]. This approach is often referred to as Dense Passage Retrieval (DPR). Similar to text, models have been proposed to encode code data [25], audio data [68], image data [24], video data [69], etc. The similarity score between dense representations are usually computed with metrics such as cosine, inner product, L2-distance.
During training, dense retrieval usually follows a contrastive learning paradigm, making positive samples more similar and negative samples less similar. Several hard negative techniques [70] have been proposed to further improve the model quality. During inference, approximate nearest neighbor (ANN) methods are applied for efficient searching. Various indices are developed to serve ANN search, such as tree [71, 72], locality sensitive hashing [73], neighbor graph index (e.g., HNSW [74], DiskANN [75]), and the combination of graph index and inverted index (e.g., SPANN [22]).
II-B3 Others
@@ III-A1 Query-based RAG
Stemming from the idea of prompt augmentation, Query-based RAG seamlessly integrates the user’s query with insights from retrieved information, feeding it directly into the initial stage of the language model’s input. This method is prevalent in RAG applications. Post-retrieval, the obtained content is merged with the user’s original query to form a composite input sequence, which is then processed by the generator to create a response. Query-based RAG is widely employed across various modalities.
For Text Generation, REALM [33] employs a dual-BERT framework to streamline knowledge retrieval and integration, marrying pre-trained models with knowledge extractors. Lewis et al. [34] leveraged DPR for information retrieval and employs BART as the generator to effectively enhance the generation. SELF-RAG [84] utilizes a critique module to determine whether the retrieval is required. In addition to being compatible with local generators, Query-based RAG is also applicable to scenarios that use LLM through API calls. REPLUG [85] follows this methodology by treating the language model as a “black box”, and effectively integrates relevant external documents into the query. In-Context RALM [86] uses BM25 for document retrieval and trains a predictive reranker to reorder and integrate the top-ranked documents.
In the field of Code, several works [87, 42, 88, 89, 90] have utilized the query-based paradigm to incorporate contextual information from text or code into the prompt, resulting in improved effectiveness of downstream tasks.
@@
Recent researches in Knowledge Base Question Answering (KBQA) has also shown significant effects of combining retrieval and language models. For instance, Uni-Parser [91], RNG-KBQA [81], and ECBRF [92] effectively improve the performance and accuracy of QA systems by merging queries and retrieved information into prompts.
@@
In the AI-for-Science field, Chat-Orthopedist [93] aids shared decision-making for adolescents with idiopathic scoliosis, improving LLMs’ effectiveness and information precision by incorporating retrieved data into model prompts.
@@ III-A2 Latent Representation-based RAG
In the Text field, FiD [35] and RETRO [36] are two classic structures of Latent Representation-based RAG, with many subsequent works being modifications based on them. FiD [35] processes each retrieved paragraph and its title alongside the query through distinct encoders, then amalgamates the resulting latent representations for decoding by a single decoder to produce the final output. RETRO [36] retrieves relevant information for each segmented sub-query, then applies a novel module termed Chunked Cross-Attention (CCA) to integrate the retrieved contents with each sub-query tokens. In addition, there are other noteworthy novel structures within the scope of latent representation-based RAG. Several studies [31, 95] have integrated kNN search within transformer blocks, allowing for input chunking and, in theory, addressing the long-criticized context length constraints of Transformer models. Kuratov et al. [96] integrated Transformer with RNN, utilizing the model’s intermediate output as the content for retrieval.
In the realms of Code, and Science, FiD has gained widespread adoption, with applications spanning code-related fields [97, 98, 99, 100, 101], and AI-for-Science [55], among others.
In the Image domain, several studies [102, 103, 104, 105] employ cross-attention mechanisms to fuse retrieval results by integrating their latent representations. Conversely, Li et al. [106] implement a text-image Affine Combination Module (ACM) that directly concatenates hidden features.
Within the Knowledge domain, several studies [107, 108, 109, 110, 111] have adopted FiD and its derivatives for downstream tasks. EaE [112] enhances the generator’s understanding through entity-specific parameterization, while TOME [113] pivots to a nuanced encoding of mentions, prioritizing the granularity of mentions over entity representations alone.
@@
For video captioning, R-ConvED [48] uses a convolutional encoder-decoder network to process retrieved video-sentence pairs with an attention mechanism, generating hidden states to produce captions. CARE [116] introduces a concept detector to produce concept probabilities, and incorporates concept representations into a hybrid attention mechanism. EgoInstructor [49] uses gated-cross attention to merge text and video features, improving the relevance and coherence of captions for egocentric videos.
Latent Representation-based RAG is versatile across modalities and tasks, merging retriever and generator hidden states, though it needs extra training for latent space alignment. This approach allows for the creation of complex, innovative algorithms that efficiently integrate retrieved information.